# On the Pitfalls of Batch Normalization for End-to-End Video Learning: A Study on Surgical Workflow Analysis

### v4

@@ 1 Introduction
Batch Normalization (BatchNorm/BN) (Ioffe and Szegedy, 2015) is a highly effective regularizer in visual recognition tasks and is ubiquitous in modern Convolutional Neural Networks (CNNs) (He et al., 2016; Szegedy et al., 2016; Tan and Le, 2019). It is, however, also considered a source for silent performance drops and bugs due to its unique property of depending on other samples in the batch and the assumptions tied to this (Brock et al., 2021; Wu and He, 2018; Wu and Johnson, 2021). Most notably, BatchNorm assumes that batches are a good approximation of the training data and only performs well when batches are large enough and sampled i.i.d. (Ioffe and Szegedy, 2015; Wu and Johnson, 2021).
@@
This generally does not hold for sequence learning, where batches contain highly correlated, sequential samples, and has led to the use of alternatives such as LayerNorm (LN) (Ba et al., 2016) in NLP (Shen et al., 2020; Vaswani et al., 2017). In video learning, BN has been studied less (Cai et al., 2021; Wu and He, 2018), despite the use of BN-based CNNs.
@@
In natural-video tasks, CNNs are only used to extract image- or clip-wise features using pretrained CNNs (e.g. (Carreira and Zisserman, 2017)) off-the-shelf. Only the temporal model, which typically does not contain BN, is trained to aggregate features over time (Abu Farha et al., 2020; Farha and Gall, 2019; Huang et al., 2020; Ishikawa et al., 2021; Ke et al., 2019; Sener et al., 2020; Wang et al., 2020; Yi et al., 2021). However, in specialized small-data domains such as surgical video, well-pretrained CNNs may not be available (Czempiel et al., 2022; Zhang et al., 2022), requiring CNNs to be finetuned, either through 2-stage (Czempiel et al., 2020) or end-to-end (E2E) training (Jin et al., 2017). The latter seems preferable to enable joint learning of visual and temporal features, especially since spatio-temporal feature extractors (e.g. 3D CNNs) have not been effective on small-scale surgical datasets (Czempiel et al., 2022; Zhang et al., 2022). However, BN layers in CNNs pose obstacles for end-to-end learning.
@@
We hypothesize that BN’s problems with correlated, sequential samples have silently caused research in video-based surgical workflow analysis (SWA) to head into a sub-optimal direction. The focus has shifted towards developing sophisticated temporal models to operate on extracted image features similar to the natural-video domain, replacing end-to-end learning with complex multi-stage training procedures, where each component (CNN, LSTM (Hochreiter and Schmidhuber, 1997), TCN (Farha and Gall, 2019), Transformer (Vaswani et al., 2017), etc.) is trained individually (Bano et al., 2020; Gao et al., 2021; Kannan et al., 2019; Marafioti et al., 2021; Yuan et al., 2021). We argue that even simple CNN-LSTM models can often outperform these methods when BN-free backbones are used and the model is trained end to end.
@@
We investigate BatchNorm’s pitfalls in end-to-end learning on two surgical workflow analysis (SWA) tasks: surgical phase recognition (Twinanda et al., 2016) and instrument anticipation (Rivoir et al., 2020). We choose these for two reasons: The lack of well-pretrained feature extractors and ineffectiveness of 3D CNNs signify the need for end-to-end approaches in SWA and thus make BN-issues most relevant here. Further, SWA is one of the most active research areas for online video understanding, where models are constrained to access only past frames, causing additional BN problems regarding leakage of future information.
@@ 2.1 Batch Normalization and Beyond
BatchNorm (Ioffe and Szegedy, 2015) is used in almost all modern CNNs. It is assumed to have a regularizing effect, improve convergence speed and enable training of deeper networks (Bjorck et al., 2018; De and Smith, 2020; Hoffer et al., 2017). In BatchNorm, features $x$ are standardized by estimating the channel-wise mean and variance and then transformed using an affine function with learnable parameters $\gamma,\beta$. Formally,
@@
Despite BN’s popularity and success, it has major disadvantages. The discrepancy between train and test behavior can lead to drops in performance when batch statistics are poor estimates of global statistics, e.g. when batches are small or contain correlated samples. Also, BN’s property of depending on other samples in the batch can lead to hidden mistakes if not handled with care. Resulting problems have been observed in many areas such as object detection (Wu and He, 2018; Wu and Johnson, 2021), sequence modeling (Ba et al., 2016), reinforcement learning (Salimans and Kingma, 2016), continual learning (Pham et al., 2022), federated learning (Andreux et al., 2020) and contrastive learning (Henaff, 2020). E.g. in language modeling, both aforementioned issues are highly problematic and have led to the use of alternative normalization layers (Ba et al., 2016; Shen et al., 2020; Vaswani et al., 2017).
Cross-Iteration BatchNorm (CBN) (Yao et al., 2021b), EvalNorm (Singh and Shrivastava, 2019) and others (Ioffe, 2017; Yan et al., 2020; Yang et al., 2021) are extensions of BN which address some of these issues. Normalization layers which only operate on single samples (e.g. GroupNorm (Wu and He, 2018), LayerNorm (Ba et al., 2016)) have also been proposed (Ba et al., 2016; Labatie et al., 2021; Liu et al., 2020; Singh and Krishnan, 2020; Ulyanov et al., 2016; Wu and He, 2018). Dynamic Normalization and Relay (DNR) (Cai et al., 2021) proposed recurrent estimation of affine parameters for video models but does not address the discrepancy of batch and global statistics.
Wu and Johnson (2021) provide the most extensive review of BN issues for vision to date, including train-test discrepancies, the effect of batch sizes and correlated samples as well as solutions such as FrozenBN (Wu and Johnson, 2021) or redefining over which samples to normalize. Related to our work, they show problems of small batches for estimating batch statistics and discuss cheating behavior in object detection and contrastive learning. Consequences for video are only hypothesized. Further, Wu and He (2018) show how BN hides benefits of higher frame rates for clip-based action classification.
Recently, there has been a trend towards CNNs without BN. Brock et al. (2021) discuss disadvantages of BN and propose a class of normalizer-free networks NFNet. They, however, require custom optimizers and are sensitive to hyperparameters. Finally, ConvNeXt (Liu et al., 2022b) is a CNN competitive with Vision Transformers (Dosovitskiy et al., 2020). It replaces BatchNorm with LayerNorm although this was not the focus of the paper.
@@ 2.2 Online Surgical Workflow Analysis
Surgical Workflow Analysis (SWA) (Maier-Hein et al., 2017) is a broad set of tasks which aim at understanding events and their interrelations in recorded or live surgical videos (Fig. 1). SWA covers a wide range of common computer vision problems including temporal action segmentation (e.g. surgical phase recognition (Twinanda et al., 2016)), dense frame-wise regression (e.g. anticipating instrument usage (Rivoir et al., 2020) or procedure duration (Twinanda et al., 2018)) or video classification (e.g. early surgery type recognition (Kannan et al., 2019)). SWA tasks are inherently video-based as they require understanding the sequence of events which have occurred until the current time point and often require memorizing long-range dependencies. The overall goal is to provide live, context-aware assistance during surgery (Maier-Hein et al., 2022). Thus, tasks are mostly defined as online recognition problems (Czempiel et al., 2020; Kannan et al., 2019; Rivoir et al., 2020; Twinanda et al., 2018).
@@
E2E without BN: Early methods completely avoid BN-issues by using backbones without BatchNorm. End-to-end CNN-LSTM models with AlexNet (Krizhevsky et al., 2012) or VGG (Simonyan and Zisserman, 2014) backbones have been proposed for phase recognition (Bodenstedt et al., 2019a; Yengera et al., 2018), anticipation (Rivoir et al., 2020), duration prediction (Rivoir et al., 2019; Yengera et al., 2018), surgery type prediction (Kannan et al., 2019) and occlusion detection (Bano et al., 2020). All methods train long sequences of several minutes. However, results are not competitive due to the use of outdated backbones.
@@
E2E with BN: SV-RCNet (Jin et al., 2017) and MTRCNet (Jin et al., 2020) are two examples of end-to-end ResNet-LSTM models for phase recognition. However, to minimize the discrepancy between batch and global statistics, training sequences are only 10 frames long (2 and 10 seconds, respectively) and processed in large batches of 100 sequences. TMRNet (Jin et al., 2021) and LRTD (Shi et al., 2020) extend SV-RCNet with attention blocks to learn long-range dependencies. Still however, sequences of only 10 frames are trained end to end. CataNet (Marafioti et al., 2021) proposes a complex, 4-stage learning process for ResNet-LSTMs to predict surgery duration. In one of the stages, end-to-end learning on long sequences is made possible by using FrozenBN with global statistics from a previous stage.
@@
Fig. 1: Two surgical workflow tasks on the Cholec80 dataset (Twinanda et al., 2016): phase recognition, a special case of temporal action segmentation, and instrument anticipation, defined as predicting the time until occurrence of an instrument within a specified horizon.
@@
Multi-stage: Recently, methods have shifted towards more complex temporal models on top of pretrained visual features. BN-related issues are circumvented by first training a CNN on randomly sampled image batches, followed by the temporal model trained on frozen image features. Methods in this style have been proposed for phase recognition (Czempiel et al., 2020, 2021; Gao et al., 2021; Yi et al., 2022; Zisimopoulos et al., 2018), duration prediction (Aksamentov et al., 2017; Bodenstedt et al., 2019b; Twinanda et al., 2018), tracking (Nwoye et al., 2019) or anticipation (Yuan et al., 2021). Most notably, TeCNO (Czempiel et al., 2020), an MS-TCN (Farha and Gall, 2019) trained on ResNet features, is the popular approach for 2-stage learning and Trans-SVNet (Gao et al., 2021), a 3-stage method which trains a Transformer on TeCNO features, is often considered state of the art among methods with CNN backbones. More recently, few multi-stage methods have achieved strong performance by using Transformer-based backbones (Chen et al., 2022; He et al., 2022; Liu et al., 2023a, b).
@@
Awareness for BN issues: Although BatchNorm has possibly influenced the move towards complex, multi-stage or short-range models, its issues have rarely been discussed. In their end-to-end model, Yengera et al. (2018) justify the use of the BN-free AlexNet with efficiency reasons. Kannan et al. (2019) simply state that their end-to-end model with VGG outperformed ResNet, but not whether this may be due the absence to BatchNorm. The authors of SV-RCNet (Jin et al., 2017) do not explain their choice of short sequences despite large batch sizes. Nwoye et al. (2019) justify their 2-stage approach through “fair comparison”. CataNet (Marafioti et al., 2021) uses FrozenBN in an end-to-end learning stage  [^1] https://github.com/aimi-lab/catanet/blob/main/train_rnn.py#L179, but this is not mentioned in the paper. Rivoir et al. (2020) briefly mention a cheating phenomenon to justify the choice for an AlexNet in instrument anticipation.
@@
3D backbones: It is noticable that surgical workflow methods almost unanimously use 2D backbones as feature extractors. The specialized surgical domain requires backbones to be finetuned and several studies suggest that the small-scale public datasets available in this domain are not sufficient to train large 3D CNNs (Czempiel et al., 2022; Zhang et al., 2022). Only few works achieve good results on larger private datasets (Czempiel et al., 2022; Zhang et al., 2021). More recently, He et al. (2022) confirmed poor performance of 3D CNNs but showed promising results using the spatio-temporal Transformer architecture Video Swin (Liu et al., 2022c) on short 1-sec. sequences.
@@ 2.3 Action Segmentation & Anticipation
Training strategies: Methods avoid BatchNorm during temporal learning by almost exclusively relying on pretrained visual features (Carreira and Zisserman, 2017) in both action segmentation (Farha and Gall, 2019; Huang et al., 2020; Ishikawa et al., 2021; Wang et al., 2020; Yi et al., 2021) and anticipation (Abu Farha et al., 2020; Furnari and Farinella, 2020; Gao et al., 2017; Ke et al., 2019; Sener et al., 2020). Whether this development can be attributed to computational costs, the availability of well-pretrained feature extractors or BatchNorm’s properties is not yet clear.
Sener et al. (2020) indicate limitations of pretrained visual features for action anticipation by showing a large performance gap compared to using ground-truth observed actions as input. The only end-to-end anticipation model for natural video of which we are aware (AVT, (Girdhar and Grauman, 2021)) does not use BatchNorm as it is entirely Transformer-based. Although not mentioned as motivation for this model choice, BN would likely have caused both train-test discrepancy and “cheating” since single-sequence batches are used. Recently, Liu et al. (2022a) show the effectiveness of end-to-end strategies for the related temporal action detection task. They use FrozenBN but do not discuss BatchNorm issues.
Online recognition: Online action understanding has mostly focused on temporal action detection (Eun et al., 2020; Lea et al., 2017; Xu et al., 2019; Zhao et al., 2020) or anticipation. Apart from few exceptions (Ghoddoosian et al., 2022), action segmentation is usually done offline (Farha and Gall, 2019; Huang et al., 2020; Ishikawa et al., 2021; Wang et al., 2020; Yi et al., 2021).
@@ 3.1 Proposed End-to-End Learning Strategies
Our premise is that end-to-end learning is preferable over multi-stage approaches and thus we propose a simple, intuitive strategy for training CNN-LSTM models on online surgical workflow tasks. We use the LSTM to demonstrate that even simple temporal aggregation can be effective when visual features are improved through end-to-end learning and temporal context is maximized. We also argue that recurrent, state-based models, where a memory state propagates information through time, are often well suited for online recognition tasks (Furnari and Farinella, 2020; Gao et al., 2017; Ghoddoosian et al., 2022). Especially for end-to-end learning on partial video sequences, the LSTM’s hidden state can be utilized to increase temporal context. Specifically, we propose:
@@
Carry-hidden training (CHT) (optional): We select batches in temporal order. This way, the LSTM’s detached hidden state can be carried across batches to further increase temporal context during training.
@@
End-to-end learning (Jin et al., 2016, 2017), single-sequence batches (Kannan et al., 2019) and CHT (Nwoye et al., 2019) have been used individually in BN-based approaches. However, only methods with BN-free backbones such as AlexNet have been able to employ these (Bodenstedt et al., 2019a) or similar (Yengera et al., 2018) strategies in combination. This is because end-to-end training is not compatible with single-sequence batches if the backbone contains BatchNorm, and by extension with CHT as it requires single-sequence batches. The following paragraph elaborates the main problems.

### v5

@@ 1 Introduction
Batch Normalization (BatchNorm/BN) [34] is a highly effective regularizer in visual recognition tasks and is ubiquitous in modern Convolutional Neural Networks (CNNs) [27, 68, 69]. It is, however, also considered a source for silent performance drops and bugs due to its unique property of depending on other samples in the batch and the assumptions tied to this [9, 76, 77]. Most notably, BatchNorm assumes that batches are a good approximation of the training data and only performs well when batches are large enough and sampled i.i.d. [34, 77].
@@
This generally does not hold for sequence learning, where batches contain highly correlated, sequential samples, and has led to the use of alternatives such as LayerNorm (LN) [4] in NLP [61, 73]. In video learning, BN has been studied less [10, 76], despite the use of BN-based CNNs.
@@
In natural-video tasks, CNNs are only used to extract image- or clip-wise features using pretrained CNNs (e.g. [11]) off-the-shelf. Only the temporal model, which typically does not contain BN, is trained to aggregate features over time [1, 19, 32, 35, 42, 60, 74, 84]. However, in specialized small-data domains such as surgical video, well-pretrained CNNs may not be available [15, 88], requiring CNNs to be finetuned, either through 2-stage [13] or end-to-end (E2E) training [37]. The latter seems preferable to enable joint learning of visual and temporal features, especially since spatio-temporal feature extractors (e.g. 3D CNNs) have not been effective on small-scale surgical datasets [15, 88]. However, BN layers in CNNs pose obstacles for end-to-end learning.
@@
We hypothesize that BN’s problems with correlated, sequential samples have silently caused research in video-based surgical workflow analysis (SWA) to head into a sub-optimal direction. The focus has shifted towards developing sophisticated temporal models to operate on extracted image features similar to the natural-video domain, replacing end-to-end learning with complex multi-stage training procedures, where each component (CNN, LSTM [30], TCN [19], Transformer [73], etc.) is trained individually [5, 24, 41, 54, 86]. We argue that even simple CNN-LSTM models can often outperform these methods when BN-free backbones are used and the model is trained end to end.
@@
We investigate BatchNorm’s pitfalls in end-to-end learning on two surgical workflow analysis (SWA) tasks: surgical phase recognition [70] and instrument anticipation [58]. We choose these for two reasons: The lack of well-pretrained feature extractors and ineffectiveness of 3D CNNs signify the need for end-to-end approaches in SWA and thus make BN-issues most relevant here. Further, SWA is one of the most active research areas for online video understanding, where models are constrained to access only past frames, causing additional BN problems regarding leakage of future information.
@@ 2.1 Batch Normalization and Beyond
BatchNorm [34] is used in almost all modern CNNs. It is assumed to have a regularizing effect, improve convergence speed and enable training of deeper networks [6, 16, 31]. In BatchNorm, features $x$ are standardized by estimating the channel-wise mean and variance and then transformed using an affine function with learnable parameters $\gamma,\beta$. Formally,
@@
Despite BN’s popularity and success, it has major disadvantages. The discrepancy between train and test behavior can lead to drops in performance when batch statistics are poor estimates of global statistics, e.g. when batches are small or contain correlated samples. Also, BN’s property of depending on other samples in the batch can lead to hidden mistakes if not handled with care. Resulting problems have been observed in many areas such as object detection [76, 77], sequence modeling [4], reinforcement learning [59], continual learning [56], federated learning [3] and contrastive learning [29]. E.g. in language modeling, both aforementioned issues are highly problematic and have led to the use of alternative normalization layers [4, 61, 73].
Cross-Iteration BatchNorm (CBN) [82], EvalNorm [65] and others [33, 79, 80] are extensions of BN which address some of these issues. Normalization layers which only operate on single samples (e.g. GroupNorm [76], LayerNorm [4]) have also been proposed [4, 44, 46, 64, 72, 76]. Dynamic Normalization and Relay (DNR) [10] proposed recurrent estimation of affine parameters for video models but does not address the discrepancy of batch and global statistics.
Wu and Johnson [77] provide the most extensive review of BN issues for vision to date, including train-test discrepancies, the effect of batch sizes and correlated samples as well as solutions such as FrozenBN [77] or redefining over which samples to normalize. Related to our work, they show problems of small batches for estimating batch statistics and discuss cheating behavior in object detection and contrastive learning. Consequences for video are only hypothesized. Further, Wu and He [76] show how BN hides benefits of higher frame rates for clip-based action classification.
Recently, there has been a trend towards CNNs without BN. Brock et al. [9] discuss disadvantages of BN and propose a class of normalizer-free networks NFNet. They, however, require custom optimizers and are sensitive to hyperparameters. Finally, ConvNeXt [50] is a CNN competitive with Vision Transformers [17]. It replaces BatchNorm with LayerNorm although this was not the focus of the paper.
@@ 2.2 Online Surgical Workflow Analysis
Surgical Workflow Analysis (SWA) [53] is a broad set of tasks which aim at understanding events and their interrelations in recorded or live surgical videos (Fig. 1). SWA covers a wide range of common computer vision problems including temporal action segmentation (e.g. surgical phase recognition [70]), dense frame-wise regression (e.g. anticipating instrument usage [58] or procedure duration [71]) or video classification (e.g. early surgery type recognition [41]). SWA tasks are inherently video-based as they require understanding the sequence of events which have occurred until the current time point and often require memorizing long-range dependencies. The overall goal is to provide live, context-aware assistance during surgery [52]. Thus, tasks are mostly defined as online recognition problems [13, 41, 58, 71].
@@
E2E without BN: Early methods completely avoid BN-issues by using backbones without BatchNorm. End-to-end CNN-LSTM models with AlexNet [43] or VGG [63] backbones have been proposed for phase recognition [7, 83], anticipation [58], duration prediction [57, 83], surgery type prediction [41] and occlusion detection [5]. All methods train long sequences of several minutes. However, results are not competitive due to the use of outdated backbones.
@@
E2E with BN: SV-RCNet [37] and MTRCNet [38] are two examples of end-to-end ResNet-LSTM models for phase recognition. However, to minimize the discrepancy between batch and global statistics, training sequences are only 10 frames long (2 and 10 seconds, respectively) and processed in large batches of 100 sequences. TMRNet [39] and LRTD [62] extend SV-RCNet with attention blocks to learn long-range dependencies. Still however, sequences of only 10 frames are trained end to end. CataNet [54] proposes a complex, 4-stage learning process for ResNet-LSTMs to predict surgery duration. In one of the stages, end-to-end learning on long sequences is made possible by using FrozenBN with global statistics from a previous stage.
@@
Fig. 1: Two surgical workflow tasks on the Cholec80 dataset [70]: phase recognition, a special case of temporal action segmentation, and instrument anticipation, defined as predicting the time until occurrence of an instrument within a specified horizon.
@@
Multi-stage: Recently, methods have shifted towards more complex temporal models on top of pretrained visual features. BN-related issues are circumvented by first training a CNN on randomly sampled image batches, followed by the temporal model trained on frozen image features. Methods in this style have been proposed for phase recognition [13, 14, 24, 85, 91], duration prediction [2, 8, 71], tracking [55] or anticipation [86]. Most notably, TeCNO [13], an MS-TCN [19] trained on ResNet features, is the popular approach for 2-stage learning and Trans-SVNet [24], a 3-stage method which trains a Transformer on TeCNO features, is often considered state of the art among methods with CNN backbones. More recently, few multi-stage methods have achieved strong performance by using Transformer-based backbones [12, 28, 48, 49].
@@
Awareness for BN issues: Although BatchNorm has possibly influenced the move towards complex, multi-stage or short-range models, its issues have rarely been discussed. In their end-to-end model, Yengera et al. [83] justify the use of the BN-free AlexNet with efficiency reasons. Kannan et al. [41] simply state that their end-to-end model with VGG outperformed ResNet, but not whether this may be due the absence to BatchNorm. The authors of SV-RCNet [37] do not explain their choice of short sequences despite large batch sizes. Nwoye et al. [55] justify their 2-stage approach through “fair comparison”. CataNet [54] uses FrozenBN in an end-to-end learning stage  [^1] https://github.com/aimi-lab/catanet/blob/main/train_rnn.py#L179, but this is not mentioned in the paper. Rivoir et al. [58] briefly mention a cheating phenomenon to justify the choice for an AlexNet in instrument anticipation.
@@
3D backbones: It is noticable that surgical workflow methods almost unanimously use 2D backbones as feature extractors. The specialized surgical domain requires backbones to be finetuned and several studies suggest that the small-scale public datasets available in this domain are not sufficient to train large 3D CNNs [15, 88]. Only few works achieve good results on larger private datasets [15, 87]. More recently, He et al. [28] confirmed poor performance of 3D CNNs but showed promising results using the spatio-temporal Transformer architecture Video Swin [51] on short 1-sec. sequences.
@@ 2.3 Action Segmentation & Anticipation
Training strategies: Methods avoid BatchNorm during temporal learning by almost exclusively relying on pretrained visual features [11] in both action segmentation [19, 32, 35, 74, 84] and anticipation [1, 22, 23, 42, 60]. Whether this development can be attributed to computational costs, the availability of well-pretrained feature extractors or BatchNorm’s properties is not yet clear.
Sener et al. [60] indicate limitations of pretrained visual features for action anticipation by showing a large performance gap compared to using ground-truth observed actions as input. The only end-to-end anticipation model for natural video of which we are aware (AVT, [26]) does not use BatchNorm as it is entirely Transformer-based. Although not mentioned as motivation for this model choice, BN would likely have caused both train-test discrepancy and “cheating” since single-sequence batches are used. Recently, Liu et al. [47] show the effectiveness of end-to-end strategies for the related temporal action detection task. They use FrozenBN but do not discuss BatchNorm issues.
Online recognition: Online action understanding has mostly focused on temporal action detection [18, 45, 78, 89] or anticipation. Apart from few exceptions [25], action segmentation is usually done offline [19, 32, 35, 74, 84].
@@ 3.1 Proposed End-to-End Learning Strategies
Our premise is that end-to-end learning is preferable over multi-stage approaches and thus we propose a simple, intuitive strategy for training CNN-LSTM models on online surgical workflow tasks. We use the LSTM to demonstrate that even simple temporal aggregation can be effective when visual features are improved through end-to-end learning and temporal context is maximized. We also argue that recurrent, state-based models, where a memory state propagates information through time, are often well suited for online recognition tasks [22, 23, 25]. Especially for end-to-end learning on partial video sequences, the LSTM’s hidden state can be utilized to increase temporal context. Specifically, we propose:
@@
Carry-hidden training (CHT) (optional): We select non-overlapping batches in temporal order. This way, the LSTM’s detached hidden state can be carried across batches to further increase temporal context during training.
@@
End-to-end learning [36, 37], single-sequence batches [41] and CHT [55] have been used individually in BN-based approaches. However, only methods with BN-free backbones such as AlexNet have been able to employ these [7] or similar [83] strategies in combination. This is because end-to-end training is not compatible with single-sequence batches if the backbone contains BatchNorm, and by extension with CHT as it requires single-sequence batches. The following paragraph elaborates the main problems.